OpenStudy (anonymous):

2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

Expand the values first 2(x+6)-(x+1)=2x+12-x-1=x+11 so x+11=16 then x=5

OpenStudy (anonymous):

Distribute first. If you know how to.

OpenStudy (anonymous):

2(x+6) becomes 2x+12 -(x+1) becomes -1(x+1) which becomes -x-1 Put it together to get 2x+12-x-1 Combine like terms: 2x-x 12-1 to get x+11 x+11=16 Solve for x.